When Red Cubs played against the Archers,15 000 seats were sold.A reserved seat cost 500.00, but admission to the unreserved seats was 350.00. The total receipts were 6 000 000.00. How many reserved seats were sold?
Answer by lwsshak3(11628)

When Red Cubs played against the Archers,15 000 seats were sold.A reserved seat cost 500.00, but admission to the unreserved seats was 350.00. The total receipts were 6 000 000.00. How many reserved seats were sold?
**
let x=number of reserved seats sold
15000-x=number of unreserved seats sold
..
500x+350*15000=6*10^6
500x+350*15000=6*10^6-1.5*3.5*10^6
500x=.75*10^6=750*10^3
x=(750/500)*10^3=1.5*10^3=1500
ans:
number of reserved seats sold=1500